OpenStudy (anonymous):

FAN AND MEDAL.!! Write the vertex form of the equation of the parabola which has a vertex of (0, 0) and contains the point (1, 2).

OpenStudy (anonymous):

\[A. y=2x ^{2} B. y=2(x-1)^{2} C. y=x ^{2} D. y=(x+1)^{2}+1\]